# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### v2

@@ V Average-case hardness implies classical simulation hardness
$||\mathcal{D}_{C}^{\prime}-\mathcal{D}_{C}||\leq\beta,$ (3)
where $\mathcal{D}_{C}$ is the ideal output distribution of the circuit $C$ and |||||$\cdot$||| | ⋅ | | represents total variation distance.
Given the total variation distance error, the above approximate sampler can have an arbitrarily large additive error for a fixed output probability. Nevertheless, it still has a comparably small additive error for most of the output probabilities due to Markov’s inequality. Accordingly, finding the average-case solution of the output probability of the ideal sampler over randomly chosen collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, up to a certain additive imprecision, is in complexity class $\rm{BPP}^{\rm{NP}}$ by Stockmeyer’s theorem [47].
Lemma 2 (Average-case approximation [1]).
@@
We leave the proof of Lemma 2 in Appendix F for a more self-contained analysis. The complexity $\rm{BPP}^{\rm{NP}}$ is known to be inside the finite level of PH, i.e., $\rm{BPP}^{\rm{NP}}\subseteq PH$. Also, by Toda’s theorem [48], PH problems can be solved given the ability to solve any #P problem, i.e., $\rm{BPP}^{\rm{NP}}\subseteq\rm{PH}\subseteq\rm{P}^{\rm{\#P}}$. If finding the average-case solution of output probabilities of sampler $\mathcal{S}$ is #P-hard, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Therefore, if an efficient classical algorithm exists that can simulate $\mathcal{S}$, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}}$ which implies the collapse of PH. Consequently, under the assumption of the non-collapse of the PH, there is no efficient classical algorithm capable of simulating $\mathcal{S}$.
@@
Suppose that the allowed additive imprecision for the problem in Theorem 4 to be #$\rm{P}$-hard can be improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$. Then, in the dilute regime such that $\gamma\geq 2$, the efficient classical simulation of approximate boson sampler $\mathcal{S}$ with respect to circuits from the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 4}$ implies the collapse of $\rm{PH}$.
@@
We lastly remark that, because our average-case hardness result in Sec. IV solely relies on collision-free outcomes, it must be ensured that these collision-free outcomes occupy a sufficient portion of all possible outcomes of Boson Sampling (for more details, see Appendix B). Hence, we focus on the dilute regime $\gamma\geq 2$ for the classical hardness result in this section, to ensure the dominance of collision-free outcomes. To similarly obtain the classical hardness result in the saturated regime $21\leq\gamma<$ 2, extending our average-case hardness result for the collision outcomes is crucial.
VI Classical simulation hardness of shallow-depth Gaussian Boson Sampling
In this section, we show that our hardness results of the shallow-depth Boson Sampling can be generalized to the Gaussian Boson Sampling scheme  [2]. Our specific setup for the Gaussian Boson Sampling is as follows. Let the total mode number $M$ of the circuit be a power of 2, and now the input state is an $M$ product of single-mode squeezed vacuum (SMSV) state |$\ket{\text{SMSV}}^{\otimes M}$ with equal squeezing parameter $r$ and equal squeezing direction. Also, let us define the output mean photon number as an integer $N$ (i.e., $N=M\sinh^{2}r$) where $M$ and $N$ are polynomially related as $M=c_{1}N^{\gamma}$ for a constant $c_{1}$ and $\gamma\geq 2$. We define $q_{\bm{s}}(C)$ as an output probability of the Gaussian Boson Sampling, for an $N$-photon outcome $\bm{s}$ from an $M$-mode linear optical circuit matrix $C$ on input $M$ SMSV states. For collision-free outcome $\bm{s}$, $q_{\bm{s}}(C)$ can be expressed as [2]
$q_{\bm{s}}(C)=\left|\bra{\bm{s}}\hat{\mathcal{U}}(C)\ket{\text{SMSV}}^{\otimes M}\right|^{2}=\frac{\tanh^{N}r}{\cosh^{M}r}|\text{Haf}((CC^{T})_{\bm{s}})|^{2},$ (6)
where |$\ket{\bm{s}}$ is an $M$-mode Fock-state corresponding to the outcome $\bm{s}$, $\hat{\mathcal{U}}(C)$ is a unitary operator corresponding to the circuit $C$, and $(CC^{T})_{\bm{s}}$ is an $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and column of the matrix $CC^{T}$.
@@ VI Classical simulation hardness of shallow-depth Gaussian Boson Sampling
Approximating the output probability $q_{\bm{s}_{0}}(C)$ of Gaussian Boson Sampling to within additive error $2^{-\frac{\gamma-1}{2}N\log N-O(N)}$ for any $C$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is #$\rm{P}$-hard in the worst case.
@@
Our strategy is to employ the scheme in Ref. [49], which used $M_{0}$ product of equally squeezed two-mode squeezed vacuum (TMSV) state as an input state to perform $M_{0}$ mode Boson Sampling task. Specifically, a single TMSV state with squeezing parameter $r$ can be represented as
|$\ket{\text{TMSV}}=\frac{1}{\cosh r}\sum_{n=0}^{\infty}\tanh^{n}r\ket{n}\ket{n},$ (7)
@@
|$\begin{split}\ket{\text{TMSV}}^{\otimes M_{0}}&=\frac{1}{\cosh^{M_{0}}r}\left(\sum_{n=0}^{\infty}\tanh^{n}r\ket{n}_{(1)}\ket{n}_{(2)}\right)^{\otimes M_{0}}\\ &=\frac{1}{\cosh^{M_{0}}r}\sum_{n=0}^{\infty}\tanh^{n}r\sum_{\bm{s}_{n}}\ket{\bm{s}_{n}}_{(1)}\ket{\bm{s}_{n}}_{(2)},\end{split}$ (8)
@@
For each mode in the given $M_{0}$ mode circuit $C_{0}$, one-half of the TMSV state (i.e., subscript (2) in Eq. (8)) is input into it, and the other half of each state (i.e., subscript (1) in Eq. (8)) is sent directly to a photon counter. By setting each |$\ket{\bm{s}_{\text{in}}}$ and |$\ket{\bm{s}_{\text{out}}}$ as a total $M_{0}$ mode and total $N_{0}$ photon Fock-state, the output probability can be represented as
$\left|\bra{\bm{s}_{\text{in}}}_{(1)}\bra{\bm{s}_{\text{out}}}_{(2)}\hat{\mathcal{U}}_{(2)}(C_{0})\ket{\text{TMSV}}^{\otimes M_{0}}\right|^{2}=\frac{\tanh^{2N_{0}}r}{\cosh^{2M_{0}}r}\left|\bra{\bm{s}_{\text{in}}}\hat{\mathcal{U}}(C_{0})\ket{\bm{s}_{\text{out}}}\right|^{2},$ (9)
@@
Note that two $M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture can be embedded in the middle of an $M=2M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture. Accordingly, we define a circuit $C$ in an $M$-mode $\mathcal{B}\mathcal{B}^{*}$ by embedding the given $M_{0}$ mode circuit $C_{0}$ in one side of $\mathcal{B}\mathcal{B}^{*}$, setting gates located right in front of the input ports as balanced beam splitters, and setting the remaining gates as identity gates; we leave in Fig. 3 an illustration of $M=16$ mode circuit $C$ for more clarity. Here, the input $M$ SMSV states with squeezing parameter $r$ combined with the balanced beam splitters at the front becomes $M_{0}$ TMSV states with squeezing parameter $r$. Therefore, our overall setup exactly follows the scheme in Ref. [49], such that the first $M_{0}$ mode is the photon counter sector to determine the input configuration of Boson Sampling, and the last $M_{0}$ mode is to simulate the Boson Sampling for the given circuit $C_{0}$.
@@
We also define an $M$-dimensional vector $\bm{s}$ as a serial concatenation of two $\bm{s}_{0}$ vectors, so that $\bm{s}$ represents an $N$-photon outcome for $M$ modes. Then the output probabilities $p_{\bm{s}_{0}}(C_{0})$ and $q_{\bm{s}}(C)$ are related as
$\begin{split}q_{\bm{s}}(C)&=\left|\bra{\bm{s}}\hat{\mathcal{U}}(C)\ket{\text{SMSV}}^{\otimes M}\right|^{2}\\ &=\left|\bra{\bm{s}_{0}}_{(1)}\bra{\bm{s}_{0}}_{(2)}\hat{\mathcal{U}}_{(2)}(C_{0})\ket{\text{TMSV}}^{\otimes M_{0}}\right|^{2}\\ &=\frac{\tanh^{2N_{0}}r}{\cosh^{2M_{0}}r}p_{\bm{s}_{0}}(C_{0}).\end{split}$ (10)
@@
Using the results of Theorem 6 and the previous proof of the average-case hardness of Boson Sampling in Theorem 4, it is straightforward to find the average-case hardness of Gaussian Boson Sampling, for a randomly chosen $N$-photon outcomes $\bm{s}\sim\mathcal{G}_{M,N}$ and a randomly chosen circuit $U\sim\mathcal{H}_{\mathcal{A}}$ in shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
@@
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $q_{\bm{s}}(U)$ of Gaussian Boson Sampling within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
It is worth emphasizing that for our setup of the Gaussian Boson Sampling, the bosonic birthday paradox is already guaranteed in the dilute regime $M=\Omega(N^{2})$, in contrast to the Boson Sampling case. More specifically, the bosonic birthday paradox requires the symmetry of inputs over the unitary circuits, such that the input pattern should be invariant under the permutations (see Appendix B for more details). Here, note that because we are considering “full mode” input, our setup already satisfies this required input symmetry. To see this symmetry clearly, one can easily check that for any unitary circuit $U$ and permutation matrix $\bm{P}$, the output probability of Gaussian Boson Sampling for $U\bm{P}$ is equivalent to
@@
Hence, we do not necessarily have to include the random permutation part $\mathcal{P}$ for the average-case hardness result; only the local random circuit for $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ suffices.
@@
The procedure is essentially the same as the proof of Theorem 4, namely, establishing a worst-to-average-case reduction from the problem in Theorem 6 to the problem in Theorem 7: First, given worst-case circuit $C\in(\mathcal{B}\mathcal{B})^{q\geq 1}$ and outcome $\bm{s}_{0}$ depicted in Theorem 6, sample permutation $\bm{P}$, such that the revised worst-case circuit and outcome become $C_{p}=\bm{P}C\in(\mathcal{B}\mathcal{B})^{q\geq 2}$ and $\bm{s}=\bm{P}\bm{s}_{0}\sim\mathcal{G}_{M,N}$ respectively. Then, we sample $V(\theta)$ from local random circuit ensemble with circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B})^{q\geq 2}$ and perturb by $C_{p}=\bm{P}C$ via Cayley transform, such that $V(\theta)$ follows the distribution $\mathcal{H}_{\mathcal{A},\theta}^{C_{p}}$. Given access to the oracle that solves the problem in Theorem 7, we infer the output probability $q_{\bm{s}}(V(\theta))$ using different values of $\theta$, and thus obtain the worst-case output probability $q_{\bm{s}}(V(1))=q_{\bm{s}}(C_{p})=q_{\bm{s}_{0}}(C)$ (for more details, see Appendix E, where we prove the average-case hardness for the Boson Sampling case).
@@
where $\mu$ is along all possible perfect matching permutations over $N$ modes. From the proof of Lemma 9, using reduction to the common denominator for all of the $m=qM\log M$ gates, $\left[V(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$. Using this fact, one can easily check that $\prod_{j=1}^{N/2}\left[(V(\theta)V(\theta)^{T})_{\bm{s}}\right]_{\mu(2j-1),\mu(2j)}$ can be represented as $(2mN,2mN)$ rational function in $\theta$, with the common denominator $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$ which does not change with $\mu$. Therefore, the output probability can be represented as $q_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with each $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ and $P(\theta)$ a degree $4mN$ polynomial function in $\theta$.
@@
VII Extension of hardness results for lossy environments
[…]
To proceed, we start with a brief review of the results presented by Ref. [50], which shows the hardness of simulating noisy quantum circuits. Specifically, one can simulate a noiseless circuit using a larger noisy circuit up to the desired imprecision, by establishing error-detecting code in the noisy circuit and post-selecting null syndrome measurements. Therefore, given the probability to post-select the no-error syndromes, one can approximate the output probability of the noiseless circuit from the output probability of the noisy circuit. Based on this argument, Ref. [28] demonstrates the average-case hardness of approximating output probabilities of noisy quantum circuits, under some plausible assumptions of the noise model. This result gives evidence of the approximate simulation hardness of noisy quantum circuits, within total variation distance error.
[…]

### v3

@@ 5 Classical simulation hardness of shallow-depth boson sampling
$\mathcal{D}_{C}^{\prime}-\mathcal{D}_{C}||\leq\beta,$ (4)
where $\mathcal{D}_{C}$ is the ideal output distribution of the circuit $C$ and $||\cdot||$ represents total variation distance.
Given the total variation distance error, the above approximate sampler can have an arbitrarily large additive error for a fixed output probability. Nevertheless, it still has a comparably small additive error for most of the output probabilities due to Markov’s inequality. Accordingly, finding the average-case solution of the output probability of the ideal sampler over randomly chosen collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, up to a certain additive imprecision, is in complexity class $\rm{BPP}^{\rm{NP}}$ by Stockmeyer’s theorem [stockmeyer1985approximation].
Lemma 3 (Average-case approximation [aaronson2011computational]).
@@
We leave the proof of Lemma 3 in Appendix G for a more self-contained analysis. The complexity $\rm{BPP}^{\rm{NP}}$ is known to be inside the finite level of PH, i.e., $\rm{BPP}^{\rm{NP}}\subseteq PH$. Also, by Toda’s theorem [toda1991pp], PH problems can be solved given the ability to solve any #P problem, i.e., $\rm{BPP}^{\rm{NP}}\subseteq\rm{PH}\subseteq\rm{P}^{\rm{\#P}}$. If finding the average-case solution of output probabilities of sampler $\mathcal{S}$ is #P-hard, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Therefore, if an efficient classical algorithm exists that can simulate $\mathcal{S}$, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}}$ which implies the collapse of PH. Consequently, under the assumption of the non-collapse of the PH, there is no efficient classical algorithm capable of simulating $\mathcal{S}$.
@@
Suppose that the allowed additive imprecision for the problem in Theorem 4 to be $\rm{P}$-hard can be improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$. Then, in the dilute regime such that $\gamma\geq 2$, the efficient classical simulation of approximate boson sampler $\mathcal{S}$ with respect to circuits from the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 4}$ implies the collapse of $\rm{PH}$.
@@
We lastly remark that, because our average-case hardness result in Sec. 4 solely relies on collision-free outcomes, it must be ensured that these collision-free outcomes occupy a sufficient portion of all possible outcomes of Boson Sampling (for more details, see Appendix B). Hence, we focus on the dilute regime $\gamma\geq 2$ for the classical hardness result in this section, to ensure the dominance of collision-free outcomes. To similarly obtain the classical hardness result in the saturated regime $1\leq\gamma<2$, extending our average-case hardness result for the collision outcomes is crucial.
6 Average-case hardness for shallow-depth Gaussian Boson Sampling
[…]
@@ 6 Average-case hardness for shallow-depth Gaussian Boson Sampling
Approximating the output probability $q_{\bm{s}_{0}}(C)$ of Gaussian Boson Sampling to within additive error $2^{-\frac{\gamma-1}{2}N\log N-O(N)}$ for any $C$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is $\rm{P}$-hard in the worst case.
@@
Our strategy is to employ the scheme in Ref. [lund2014boson], which used $M_{0}$ product of equally squeezed two-mode squeezed vacuum (TMSV) state as an input state to perform $M_{0}$ mode Boson Sampling task. Specifically, a single TMSV state with squeezing parameter $r$ can be represented as
[…]
@@
[…]
@@
[…]
@@
Note that two $M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture can be embedded in the middle of an $M=2M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture. Accordingly, we define a circuit $C$ in an $M$-mode $\mathcal{B}\mathcal{B}^{*}$ by embedding the given $M_{0}$ mode circuit $C_{0}$ in one side of $\mathcal{B}\mathcal{B}^{*}$, setting gates located right in front of the input ports as balanced beam splitters, and setting the remaining gates as identity gates; we leave in Fig. 3 an illustration of $M=16$ mode circuit $C$ for more clarity. Here, the input $M$ SMSV states with squeezing parameter $r$ combined with the balanced beam splitters at the front becomes $M_{0}$ TMSV states with squeezing parameter $r$. Therefore, our overall setup exactly follows the scheme in Ref. [lund2014boson], such that the first $M_{0}$ mode is the photon counter sector to determine the input configuration of Boson Sampling, and the last $M_{0}$ mode is to simulate the Boson Sampling for the given circuit $C_{0}$.
@@
We also define an $M$-dimensional vector $\bm{s}$ as a serial concatenation of two $\bm{s}_{0}$ vectors, so that $\bm{s}$ represents an NN-photon outcome for $M$ modes. Then the output probabilities $p_{\bm{s}_{0}}(C_{0})$ and $q_{\bm{s}}(C)$ are related as
[…]
@@
Using the results of Theorem 6 and the previous proof of the average-case hardness of Boson Sampling in Theorem 4, it is straightforward to find the average-case hardness of Gaussian Boson Sampling, for a randomly chosen NN-photon outcomes $\bm{s}\sim\mathcal{G}_{M,N}$ and a randomly chosen circuit $U\sim\mathcal{H}_{\mathcal{A}}$ in shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
@@
The following problem is $\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $q_{\bm{s}}(U)$ of Gaussian Boson Sampling within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
It is worth emphasizing that for our setup of the Gaussian Boson Sampling, the bosonic birthday paradox is already guaranteed in the dilute regime $M=\Omega(N^{2})$, in contrast to the Boson Sampling case. More specifically, the bosonic birthday paradox requires the permutation symmetry of inputs over the unitary circuits, in the sense that the input pattern should be invariant under the permutations (see Appendix B for more details). Since we are considering “full mode" input, our setup already satisfies this required input symmetry. To see this input symmetry clearly, one can easily check that for any unitary circuit $U$ and permutation matrix $\bm{P}$, the output probability of Gaussian Boson Sampling for $U\bm{P}$ is equivalent to
@@
Therefore, we do not necessarily have to include the random permutation part $\mathcal{P}$ for the average-case hardness result; only the local random circuit for $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ suffices.
@@
The procedure is essentially the same as the proof of Theorem 4, namely, establishing a worst-to-average-case reduction from the problem in Theorem 6 to the problem in Theorem 7: First, given worst-case circuit $C\in(\mathcal{B}\mathcal{B})^{q\geq 1}$ and outcome $\bm{s}_{0}$ depicted in Theorem 6, sample permutation $\bm{P}$, such that the revised worst-case circuit and outcome become $C_{p}=\bm{P}C\in(\mathcal{B}\mathcal{B})^{q\geq 2}$ and $\bm{s}=\bm{P}\bm{s}_{0}\sim\mathcal{G}_{M,N}$ respectively. Then, we sample $V(\theta)$ from local random circuit ensemble with circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B})^{q\geq 2}$ and perturb by $C_{p}=\bm{P}C$ via Cayley transform, such that $V(\theta)$ follows the distribution $\mathcal{H}_{\mathcal{A},\theta}^{C_{p}}$. Given access to the oracle that solves the problem in Theorem 7, we infer the output probability $q_{\bm{s}}(V(\theta))$ using different values of $\theta$, and thus obtain the worst-case output probability $q_{\bm{s}}(V(1))=q_{\bm{s}}(C_{p})=q_{\bm{s}_{0}}(C)$ (for more details, see Appendix F, where we prove the average-case hardness for the Boson Sampling case).
@@
where $\mu$ is along all possible perfect matching permutations over NN modes. From the proof of Lemma 10, using reduction to the common denominator for all of the $m=qM\log M$ gates, $\left[V(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$. Using this fact, one can easily check that $\prod_{j=1}^{N/2}\left[(V(\theta)V(\theta)^{T})_{\bm{s}}\right]_{\mu(2j-1),\mu(2j)}$ can be represented as $(2mN,2mN)$ rational function in $\theta$, with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)]^{N}$ which does not change with $\mu$. Therefore, the output probability can be represented as $q_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with each $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ and $P(\theta)$ a degree $4mN$ polynomial function in $\theta$.
@@
7 Hardness analyses in lossy environments
In this section, we generalize our hardness results for $\it{lossy}$ environments, namely, shallow-depth linear optical circuits suffering from photon loss channels after each gate implementation. The reason we consider such a noise channel is that photon loss is indeed a major source of error in optical systems [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian]. Also, photon loss ruins the classical intractability of Boson Sampling, as there exist many efficient classical algorithms that can simulate lossy Boson Sampling within a constant total variation distance [oszmaniec2018classical, garcia2019simulating, qi2020regimes]. Therefore, we mainly deal with the photon loss error here; our goal is to provide evidence for the hardness of the approximate simulation of Boson Sampling in lossy shallow circuits within total variation distance error. For simplicity, we do not consider any photon gain error here, such as thermal radiation noise subjected to the circuits.
To proceed, we start with a brief review of the results presented by Ref. [fujii2016noise], which shows the hardness of simulating noisy quantum circuits. Specifically, one can simulate a noiseless circuit using a larger noisy circuit up to the desired imprecision, by establishing error-detecting code in the noisy circuit and post-selecting null syndrome measurements. Therefore, given the probability to post-select the no-error syndromes, one can approximate the output probability of the noiseless circuit from the output probability of the noisy circuit. Based on this argument, Ref. [bouland2022noise] demonstrates the average-case hardness of approximating output probabilities of noisy quantum circuits, under some plausible assumptions of the noise model. This result gives evidence of the approximate simulation hardness of noisy quantum circuits, within total variation distance error.
The main strategy of the above hardness results is approximating ideal output probabilities by post-selecting error-free results from noisy circuits. Here, we can directly apply their strategy to our case, i.e., lossy shallow-depth linear optical circuits. The crucial observation is that considering photon loss error on Boson Sampling, the error syndrome is the output photon number $\it{itself}$. Specifically, if the output photon number is the same as the input photon number, this implies that no loss occurred throughout the circuit. Therefore, by post-selecting the event that the measured output photon number is the same as the input photon number, we can infer ideal output probabilities.